Math Probability Question
Jayne has a collectionof 50 old teapots. 27 of them have rose patterns. Some of them are covered in blue willow patterns. Some of them have striped patterns and the rest have plain colours.
One teapot is chosen at random:
The probability that it will have a striped pattern is 0.08
How many of Jaynes teapots are striped? The probability that it will have a blue willow pattern is 0.22.
How many of Jaynes teapots have plain colours.

Answer by stanbon(75887)

Jayne has a collectionof 50 old teapots.
27 of them have rose patterns: 0.54 have rose patterns
Some of them are covered in blue willow patterns.
Some of them have striped patterns and the rest have plain colours.
-----------------------------------------
One teapot is chosen at random:
The probability that it will have a striped pattern is 0.08
-----------------------------------------
How many of Jaynes teapots are striped?
# of striped = 0.08*50 = 4
-----------------------------------------
The probability that it will have a blue willow pattern is 0.22.
How many of Jaynes teapots have plain colours.
% of plain = 100%-(0.54+0.08+0.22) = 0.16
# of plain colored = 0.16*50 = 8
=================================
Cheers,
Stan H.